OpenStudy (winsteria):

WILL MEDAL AND FAN What is 1 and 3/14 - (-2 and 1/7) ? Express the answer in simplest form.

OpenStudy (anonymous):

ask @leonardo0430

OpenStudy (anonymous):

i got 179/28

OpenStudy (ageta):

179/28

OpenStudy (anonymous):

which is 6.392857143

OpenStudy (ageta):

6.4

OpenStudy (winsteria):

It has to be simplified as a mixed fraction for the answer though

OpenStudy (winsteria):

@aum

OpenStudy (ageta):

Write the decimal number as a fraction (over 1) 6.39 = 6.39 / 1 Multiplying by 1 to eliminate 2 decimal places we multiply top and bottom by 2 10's Numerator (N) N = 6.39 × 10 × 10 = 639 Denominator (D) D = 1 × 10 × 10 = 100 N / D = 639 / 100 Simplifying our fraction = 639/100 = 639/100 = 6 39/100

OpenStudy (ageta):

i think this help

OpenStudy (aum):

\[ 1\frac{3}{14} - (-2\frac 17) = 1\frac{3}{14} + 2\frac 17 = 1 + 2 + \frac{3}{14} + \frac 17 = 3 + \frac{3}{14} + \frac {2}{14} = 3\frac{5}{14} \]

OpenStudy (winsteria):

Okay so the answer is 3 5/14 none of the other ones right?

OpenStudy (aum):

That is the answer I am getting. If you need further explanation on any of the steps you can ask me.

OpenStudy (winsteria):

Alright thank you so much

OpenStudy (aum):

You are welcome.